# Cardinalities in proofreading

### v1

@@ Abstract
Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the bound
${\textstyle\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
Two analogous bounds hold when $A_{Y,X}$ is replaced by $A_{Y,X}^{\prime}$, resp. $C_{Y,X}$, and $pm$ by $qm$, resp. $pqm$.
We explain in detail this restatement. Let $m\in\mathbb{N}$, $p\in[0,1]$, $X_{i}$ for $i\in[m]$ be $m$ independent two-valued real random variables in a probability space, with the common distribution $\mathrm{Pr}(X_{i}=1-p)=p$ and $\mathrm{Pr}(X_{i}=-p)=1-p$, and let
@@
The random variable $X_{0}$ has the centered binomial distribution $B(m,p)-pm$. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
@@
For every real number $a>0$,
${\textstyle\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.}$
What does the independence of random variables $X_{i}$, $i\in[m]$, which is assumed in Theorem 2.4, precisely mean? Some events $C_{1}$, $C_{2}$, $\dots$, $C_{m}$ in a probability space are independent, if for every set $I\subset[m]$ we have
${\textstyle\mathrm{Pr}\big{(}\bigcap_{i\in I}C_{i}\big{)}=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.}$
These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent. A well known lemma, whose proof we leave for the interested reader as an exercise, says that any independent events are strongly independent. Discrete (i.e., with at most countable images) real random variables $Y_{1}$, $Y_{2}$, $\dots$, $Y_{m}$ are independent if for every $m$-tuple
$\overline{a}=(a_{1},\,a_{2},\,\dots,\,a_{m})\in\mathbb{R}^{m}$
the events $\mathrm{Pr}(Y_{i}=a_{i})$, $i\in[m]$, are independent. If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $c_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
We deduce Theorem 2.3 from Theorem 2.4. We consider only events $A_{Y,X}$, for events $A_{Y,X}^{\prime}$ and $C_{Y,X}$ we would argue similarly, relying on part 2 and 3 of Proposition 2.2. So let $(P,\overline{S})$, with
@@
be as above two $(p,q,n)$-proofreaders and let $X\subset[n]$ with $|X|=m\geq 1$. We may assume that $X=[m]$. For $i\in[m]$ we define two-valued $P(MN)$-random variables $X_{i}$ by setting $X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$. We set $X_{0}=\sum_{i=1}^{m}X_{i}$. Then for any subset $Y\subset X$ we have
@@
By part 1 of Proposition 2.2 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent. Let $a>0$ be a real number. By applying Theorem 2.4 to the random variable $X_{0}$ we get the estimate in Theorem 2.3.
We proceed to proofs of Propositions 2.1 and 2.2.
Proof of Proposition 2.1. Suppose that $(P,\overline{S})$, where
@@
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Let $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$. We may assume that the $1$s of $\overline{t}$ are on the first $m$ positions. We consider the $P(MN)$-events $A_{i}$, $i\in[m]$. Let $Y\subset[m]=X$. Then
${\textstyle\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $s_{v_{j},i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.}$
@@
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big{\}}\big{|}=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.}$
The stated estimate therefore follows from the first part of Theorem 2.3. The second part of Proposition 2.1 follows by a very similar argument from the second part of Theorem 2.3. $\Box$
Proof of Proposition 2.2. Let $(P,\overline{S})$, $a>0$ and $\overline{t}$ be as in the previous proof. We consider the $P(MN)$-events $C_{i}$, $i\in[m]$, and argue as in the previous proof, relying on the third part of Theorem 2.3. $\Box$
@@ 3 Concluding remarks
By [5, p. 109], “The subject of probability is wholly free from real paradoxes.” This is not precise — not only books like [10] are devoted to paradoxes in probability theory, but some of these paradoxes are truly real. At their roots is the unrealistic nature, or even total absence, of physical experiments that should be modeled by the probabilistic model, or PM, at hand. In these paradoxical cases we often have a PM that does not model anything realistic. For example, take well known Bertrand’s paradox which asks about the probability that random chord in the unit circle in the plane has length at least $\sqrt{3}$ (the side of the inscribed equilateral triangle). The noun phrase “random chord” is problematic because it does not refer to anything clearly defined. In contrast, noun phrases like “random (finite!) graph” or “random subset (of a finite set)” are unproblematic; these concepts are easily modeled by finite uniform probability spaces and events in them, like the above $P(MN)$ and the $P(MN)$-events $A_{i}$, $A_{i}^{\prime}$ and $C_{i}$, $i\in[n]$. Mathematical chords have zero width and form an uncountable set. People try to resolve Bertrand’s paradox by rushing in defining PM on the set of mathematical chords. Instead, they should step back and consider physical chords with width $w>0$. A realistic and convincing physical discrete model, or DM, for selection of random physical chords should be devised. Only when such DM is understood, we can send $w$ to 00 and see what happens. This approach is implemented in [6]. We hope to rework [6] in near future in a better form.
In this article we approached Pólya’s proofreaders from the DM perspective. Returning to Proposition 1.1, we see that if $(P,\overline{s})$, where
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,,$
is the $(p,n)$-proofreader, then in the uniform probability space
$P(N)=([N],\,\mathcal{P}([N]),\,\mathrm{Pr})\,,$
with $\mathrm{Pr}(\{j\})=\frac{1}{N}$, the $n$ $P(N)$-events
$A_{i}=\{j\in[N]\colon\;s_{j,i}=1\},\ i\in[n]\,,$
are independent and model random subset of $[n]$ with probability of an element equal to $p$. However, our construction requires $p$ ($\in(0,1]$) be a rational number. This is just as right because we cannot have a realistic physical experiment with probabilities like $\frac{1}{\sqrt{2}}$. Thus we conclude our article with the comment, which may come as a surprise to some, that the random graph $G(n,p)$ with probability of an edge equal to an irrational number $p$ is not really a realistic model.

### v2

@@ 3. Now the displayed product equals
Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the following bounds.
  1. 1.
    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  2. 2.
    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-qm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X}^{\prime})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
  3. 3.
    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pqm\,|\leq a\end{subarray}}\mathrm{Pr}(C_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
We explain in detail the deduction of this theorem from [2, Corollary A.1.7]. Let $m\in\mathbb{N}$, $p\in[0,1]$, $X_{i}$ for $i\in[m]$ be $m$ independent two-valued real random variables in a probability space, with the common distribution $\mathrm{Pr}(X_{i}=1-p)=p$ and $\mathrm{Pr}(X_{i}=-p)=1-p$, and let
@@
The random variable $X_{0}$ has the centered binomial distribution B(m,p)pmB(m,p)-pm. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
@@
With the above notation we have for every real number $a>0$ that
$\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.$}
What does the independence of random variables $X_{i}$, $i\in[m]$, which is needed in Theorem 2.3, precisely mean? Some events $C_{1}$, $C_{2}$, $\dots$, $C_{m}$ in a probability space are independent, if for every set $I\subset[m]$ we have
$\mathrm{Pr}\big(\bigcap_{i\in I}C_{i}\big)=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.$}
These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where each $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent. A well known lemma, whose proof we leave for the interested reader as an exercise, says that any independent events are strongly independent. Discrete (i.e., with at most countable images) real random variables $Y_{1}$, $Y_{2}$, $\dots$, $Y_{m}$ are independent if for every $m$-tuple
$\overline{a}=(\overline{a}_{1},\,\overline{a}_{2},\,\dots,\,\overline{a}_{m})\in\mathbb{R}^{m}\,$,
the events $\mathrm{Pr}(Y_{i}=\overline{a}_{i})$, $i\in[m]$, are independent. If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $\overline{c}_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
We deduce Theorem 2.2 from Theorem 2.3. We consider only item 1, the arguments for items 2 and 3 are similar. So let $(P,\overline{S})$, with
@@
be as above two $(p,q,n)$-proofreaders and let $X\subset[n]$ with $|X|=m\geq 1$. We may assume that $X=[m]$. For $i\in[m]$ we define two-valued $\mathrm{PS}(MN)$-random variables $X_{i}$ by setting ($j\in[MN]$)
$\text{$X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$}\,.$
We set $X_{0}=\sum_{i=1}^{m}X_{i}$. Then for any subset $Y\subset X$ we have
@@
By item 1 of Proposition 1.4 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent. Let $a>0$ be a real number. Applying Theorem 2.3 to the random variable $X_{0}$ we get the estimate in item 1 of Theorem 2.2.
We proceed to the proof of Proposition 2.1.
Proof of Proposition 2.1. We consider only item 1, the arguments for items 2 and 3 are similar. Suppose that $(P,\overline{S})$, where
@@
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Let $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$. We may assume that the $1$s of $\overline{t}$ are on the first $m$ positions. We consider the $\mathrm{PS}(MN)$-events $A_{i}$, $i\in[m]$. Let $Y\subset[m]=X$. Then
$\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $(\overline{s}_{v_{j}})_{i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.$}
@@
$\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big\}\big|=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.$}
The stated estimate therefore follows from item 1 of Theorem 2.2. $\Box$
@@ 3 Concluding remarks
An interesting direction for further research is the following. Suppose that $(P,\overline{S})$, where
are two $(p,q,n)$-proofreaders and that $\overline{t}\in B_{n}$ is a text. We do not have access to $\overline{t}$ and to $m=\|\overline{t}\|$, but the multiset of triples of values
$V=\{(P(\overline{s}_{v_{j}},\,\overline{t}),\,P(\overline{u}_{w_{j}},\,\overline{t}),\,P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t})))\in\mathbb{N}_{0}^{3}\colon\;j\in[MN]\}$
is known. Can we recover the number of errors $m$, or at least some bounds on $m$, from $V$? The parameter $n$ is of course known but we get two versions of this problem depending on whether $p$ and $q$ are known or not.
Definitions 1.1 and 1.3 with accompanying Propositions 1.2 and 1.4 are probably more important results than Theorem 1.5 whose deduction from the well known Theorem 2.3 is relatively straightforward. The two definitions provide realistic models (RM) of Pólya’s proofreaders. Indeed, if
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})$
are runs of a $(p,n)$-proofreader $(P,\overline{s})$ and $\overline{t}\in B_{n}$ is a text, then we can, at least in principle, physically go through all $N=l^{n}$ binary $n$-tuples $\overline{s}_{j}$ and compute all numbers $\|P(\overline{s}_{j},\overline{t})\|$ of errors the proofreader detects in the text during all $N$ runs.
For this to be possible, it is crucial that $p\in\mathbb{Q}$. If $p=1/\sqrt{2}$, say, then no RM for a proofreader with efficiency $p$ exists. For irrational $p$ one has to be content with probabilistic models (PM), of random subsets containing (independently) a given element with probability $p$. Their construction is easy, but they cannot be physically realized in the same sense in which RM can.
Some paradoxes in probability theory arise when a “practical” problem is modeled right away with a PM, while one should start carefully with a RM. The best example of this is the well known Bertrand paradox. This paradox begins as a “practical” problem: what is the probability that a random chord in the unit circle is longer than the length $\sqrt{3}$ of the side of the inscribed equilateral triangle. Solutions start with various PMs, which automatically assume dimensionless points and chords with zero width. This is a mistake, no such objects exist in reality. Instead, one should start with a RM where points have size $s>0$ and chords have width $w>0$. Once such RM is established, one can send $s,w\to 0$ and see what happens in the limit. This is the approach in [5], which we hope to rework and improve in the near future.